# Defending Multimodal Backdoored Models by Repulsive Visual Prompt Tuning

### v1

@@ 4.5 Qualitative Analysis
Figure 5: Attention map for (a) original image and (b) poisoned image. The second column shows the attention map of the backdoored visual encoder attacked by BadCLIP. The third and fourth columns show the attention map of RVPT’s patch embeddings and the prompt embeddings. RVPT effectively reduces the over-attention of both the class-irrelevant features and the trigger pattern.
Attention map.  We plot the attention maps for the backdoored encoder and the recovered encoder by RVPT in Figure 5 to observe the areas that the model focuses on. Firstly, for the clean image, RVPT enables the encoder to focus more on the useful features of the bird. In contrast, the backdoored encoder also attends to irrelevant features that humans cannot interpret. Secondly, for the poisoned image, although RVPT has noticed the trigger pattern, unlike the backdoored encoder, it will not draw all the attention to it.
Figure 6: The t-SNE [47] plots for the representations of clean (dotted) and poisoned (star-shaped) images. The model is attacked by BadCLIP. We can see that RVPT successfully breaks the clustering effect of the poisoned image features in the embedding space. This clustering effect of poisoned representations was revealed in [2].
t-SNE.  We draw the t-SNE plots [47] in Figure 6 to observe the distribution of the image representations of different visual encoders. The plots show that image representations encoded by the backdoored CLIP and VPT-defended CLIP are clustered, indicating that these models focus solely on the trigger pattern, ignoring the information from the original image. On the contrary, RVPT not only disrupts this cluster but also restores the poisoned image representations, bringing them closer to their original counterparts. This observation suggests that after RVPT, the trigger pattern will have a minimal effect on the visual representations.
@@ 5 Conclusion
In this paper, we, for the first time, investigated how to utilize limited clean samples and tunable parameters to defend a backdoored multimodal contrastively pre-trained model (i.e., CLIP). We empirically showed that the excessive encoding of class-irrelevant features causes CLIP’s poor resistivity to input perturbations, resulting in its vulnerability to backdoor attacks. To reduce the class-irrelevant features, we proposed a novel method called Repulsive Visual Prompt Tuning. This method leverages our designed feature-repelling loss to minimize the mean cosine similarity between the prompted features and the original features while optimizing cross-entropy loss to maintain clean accuracy. Comprehensive experimental results on multiple attacks and benchmark datasets demonstrated the effectiveness of our proposed method.
Supplemental Material
Organization of the Supplemental Material
  • Section A provides more results of different perturbations and shows the negative correlation between PR and ASR.
@@
  • Section C shows the experiment results of combining the proposed method and other backdoor defense methods.
  • Section D contains more experiments of ablation study, such as the distance measure function, the permutation of the repulsive prompts, etc. It also contains the setting of experiments that ablates the FR loss.
  • Section E compares computational expenses between our method and CleanCLIP.
A More Analysis of Perturbation Resistivity
A.1 More Types of Perturbation
We evaluate the perturbation resistivity of four models: ViT/B32 of OpenAI CLIP, ViT supervised exclusively on ImageNet, RVPT with 1 shot on ImageNet, and RVPT with 16 shots on ImageNet against various input perturbations on the ImageNet validation set. The visual illustration is shown in Figure 1, and the results of the PR values for different models are in Table 1. The results in Table 1 show that CLIP has generally much less PR value against various perturbations than traditionally supervised ViT, and RVPT can consistently increase its PR value across all scenarios. Even with 1 shot, RVPT can effectively increase CLIP’s PR value and thus successfully defend it from various backdoor attacks.
Figure 1: Examples of the experimented perturbation. We adopt various perturbations to verify CLIP’s intrinsically low PR value.
Table 1: Perturbation resistivity of the four models across various perturbations. We can see the PR value of CLIP ViT is less than supervised ViT, while RVPT significantly improves its PR.
Perturbation ImageNet
Supervised ViT CLIP ViT VPT - 16 shot RVPT - 1 shot RVPT - 16 shot
Gaussian noise (std=0.01) 0.959 0.572 0.737 0.942 0.965
Gaussian noise (std=0.03) 0.951 0.370 0.583 0.906 0.931
Gaussian blur (kernel size=3) 0.976 0.882 0.886 0.986 0.992
Gaussian blur (kernel size=5) 0.975 0.819 0.890 0.977 0.986
Color jitter 0.946 0.832 0.887 0.974 0.983
Horizontal flipping 0.971 0.950 0.968 0.991 0.994
BadNet attack 0.976 0.861 0.954 0.982 0.991
Blended attack 0.975 0.855 0.915 0.980 0.990
BadCLIP attack (trigger pattern) 0.975 0.871 0.550 0.974 0.988
A.2 The relationship between PR and ASR
We evaluate the relationship using two kinds of perturbations, including Gaussian noise and non-adversarial trigger patterns. From Figure 2, we can see a clear, strong negative correlation between PR and ASR for different types of backdoored models.
Figure 2: The correlation between PR and ASR. By varying the prompted depth in RVPT, we recorded the PR and ASR for different depths. The result shows a clear negative correlation between PR and ASR. In (a), we use Gaussian noise whose standard deviation is 0.01; In (b), we use Blended to perturb the images fed into the CLIP infected by BadNet. For CLIP infected by BadCLIP and Blended, we use the perturbation of Random. Therefore, it justifies the motivation to reduce PR.
@@ B.1 Detailed Settings of Backdoor Attacks
In our experiments, we evaluate seven prominent backdoor attack methods: BadNet [15], Blended [6], BadCLIP [28], ISSBA [25], Label Consistent [46], WaNet [37], and TrojVQA [48]. Below, we provide a detailed description of each method:
@@
  • Label Consistent improves backdoor attack stealth from the perspective of label manipulation. It employs generative models or adversarial perturbations to selectively poison images that belong to the target class, maintaining consistency between poisoned images and their original class labels to escape human inspection. We use BadNet-LC in our experiment.
@@
  • TrojVQA proposes a dual-key backdoor attack that sets triggers on both modalities. To evaluate RVPT under backdoored generative VLMs, we use TrojVQA as the backdoor attack and use the following settings. Firstly, we optimize the visual triggers for the CLIP visual backbone with the description “This is a yellow banana.”. Specifically, We set the patch size to 16$\times$16 and locate the patch in the middle of the image. Then, we randomly corrupt the text for poisoned samples into sentences that describe the target class of banana. Moreover, we add the trigger word “SUDO ” in front of each sentence to make sure two keys for one backdoor. In the inference stage, the optimized trigger is added to the image, and “SUDO ” is added to the beginning of the prompt.
We employ the AdamW optimizer with an initial learning rate of 1e-5 for all the backdoor attack methods except 1.5e-5 for WaNet. The learning rate follows a cosine decay schedule over five epochs, and we set the batch size to 128. The visualization of each trigger pattern is in Figure 3.
@@ B.2 Detailed Settings of Backdoor Defenses
CleanCLIP  CleanCLIP [2] provides a defense against backdoor attacks in multimodal contrastive learning by optimizing the integration of multimodal contrastive and unimodal self-supervised losses, relying on a limited amount of clean data. It is important to note that CleanCLIP utilizes ResNet-50 as the backbone of its visual encoder. In contrast, our work adopts the Vision Transformer (ViT-B/32) as the visual encoder, and we have accordingly adjusted the parameters to fit this architecture. Specifically, we randomly selected 100,000 image-text pairs from the CC3M dataset for fine-tuning. The learning rates were set to 5e-6 for BadNet, TrojVQA, Blended, Label Consistent, and BadCLIP, while for WaNet and ISSBA on ImageNet-1K, they were set to 3e-6. We used a batch size of 64 and fine-tuned the models over 10 epochs. Notably, we did not focus solely on reducing the attack success rates by manipulating learning rates; instead, we ensured that the clean accuracy of the fine-tuned model was consistently maintained throughout.
@@ B.3 Detailed Settings of Datasets
C Results of Defense Combination
Table 2 shows that RVPT still shows superiority when defending against multiple backdoor attacks on ImageNet. Moreover, when RVPT is combined with a Linear Probe, the model overfits the dataset with significantly dropping accuracy. In summary, CleanCLIP + RVPT is the best choice to defend against backdoor attacks.
[…]
             Method              ImageNet
             BadNet              Blended              BadCLIP
             CP + LP              0.60 (54.64)              0.24 (56.24)              53.30 (55.79)
             RVPT + LP              0.26 (51.39)              0.04 (50.26)              0.09 (49.53)
             CP + RVPT              0.54 (55.65)              0.03 (57.00)              0.37 (56.45)
@@ D More Experiments for Ablation Study
D.1 Ablation Study of the Distance Measure Function
Instead of using cosine similarity, we experimented with maximizing the L1 and L2 distances of the normalized features, as shown in Table 3. However, both L1 and L2 losses demonstrate suboptimal performance.
[…]
           Method            ImageNet
           BadNet            blended            BadCLIP
           L1 distance            46.50 (60.20)            16.84 (60.14)            0.04 (53.72)
           L2 distance            48.72 (60.40)            77.44 (60.30)            94.73 (63.65)
           cos similarity            1.84 (58.09)            0.02 (57.97)            2.76 (61.72)
D.2 Ablation Study of the Permutation of the prompts and FR loss
To evaluate the best permutation of the prompts and FR loss, we fix the number of repulsive prompts and repelled features and design 4 kinds of permutations: Split, Overlap, Intertwine, Mix in Figure 4 and show the results in Table 4. From the result, we can conclude that repulsive prompts and repelled features should be close to the output to obtain optimal performance of both recognition and defense.
Figure 4: Visualization of the different permutations: red bars indicate that the layer’s feature is prompted, blue bars indicate that the layer’s feature is repelled, and purple bars indicate that the layer’s feature is both prompted and repelled.
[…]
               Permutation                ImageNet
               BadNet                blended                BadCLIP
               Split                7.89 (53.79)                0.03 (53.60)                12.49 (57.99)
               Overlap                1.38 (57.20)                0.05 (57.20)                7.16 (61.24)
               Intertwine                14.72 (55.47)                0.02 (55.44)                27.02 (59.98)
               Mix                9.17 (57.33)                0.02 (56.92)                9.48 (61.02)
D.3 Ablation Study of the Handcrafted Prompt
@@
           “<CLS>”            2.10 (58.05)            0.02 (58.00)            3.55 (61.73)
           “###### <CLS>”            1.91 (57.62)            0.01 (57.47)            3.76 (61.36)
           “a photo of a <CLS>”            1.84 (58.09)            0.02 (57.97)            2.76 (61.72)
D.4 Ablation Study of the Visual Encoder Architecture
@@
D.5 Settings for Experiment of Ablation Study
@@
  • For the evaluation of open-vocabulary ASR, we tune the infected model whose target class is banana with the few-shot training samples in ImageNet, which does not have the images of the target class.
@@
E Computational Expenses
We conducted the main experiments on a single NVIDIA 3090 using a half-precision data type and recorded one training process in Table 7. The results indicate that our method substantially reduces computational costs, enhancing defense efficiency.
CleanCLIP 4:46:21 13854 MB 126 M 100K
RVPT on ImageNet 45:05 3382 MB 0.34 M 16 K
RVPT on OxfordPets 2:19 1010 MB 0.34 M 0.6 K

### v2

@@ 4.5 Qualitative Analysis
Figure 5: Attention map for (a) original image and (b) poisoned image. The second column shows the attention map of the backdoored visual encoder attacked by BadCLIP. The third and fourth columns show the attention map of RVPT’s patch embeddings and the prompt embeddings. RVPT effectively reduces the over-attention of both the non-essential features and the trigger pattern.
Attention map.  We plot the attention maps for the backdoored visual encoder and the defended visual encoder by RVPT in Figure 5 to observe the areas that the model focuses on. Firstly, for the clean image, RVPT enables the encoder to focus more on the useful features of the bird. In contrast, the backdoored encoder also attends to irrelevant features that humans cannot interpret. Secondly, for the poisoned image, although RVPT has noticed the trigger pattern, unlike the backdoored encoder, it will not draw all the attention to it.
t-SNE.  We draw the t-SNE plots [51] in Figure 6 to observe the visualizations of the image representations of different visual encoders. The plots show that image representations encoded by the backdoored CLIP and VPT-defended CLIP are clustered, indicating that these models focus solely on the trigger pattern, ignoring the information from the original image. On the contrary, RVPT not only disrupts this cluster but also restores the poisoned image representations, bringing them closer to their original counterparts. This observation suggests that after RVPT, the trigger pattern will have a minimal effect on the visual representations.
Figure 6: The t-SNE [51] plots for the representations of clean (dotted) and poisoned (star-shaped) images. The model is attacked by BadCLIP [31]. RVPT successfully breaks the clustering effect of the poisoned image features in the embedding space and ensures robust representations of attacked images, as observed in [3].
@@ 5 Conclusion
In this paper, we, for the first time, investigated how to utilize limited clean samples and tunable prompts to defend a backdoored multimodal contrastively pre-trained model (i.e., CLIP). We empirically showed that CLIP’s low visual feature resistivity to input perturbations results in its vulnerability to backdoor attacks. To enhance this resistivity, we proposed a novel method called Repulsive Visual Prompt Tuning that guides CLIP to exclusively encode in-dataset predictive features. This method leverages our designed feature-repelling loss to minimize the mean cosine similarity between the prompted features and the original features across layers while optimizing cross-entropy loss to maintain clean accuracy on the downstream tasks. Comprehensive experimental results on multiple attacks and benchmark datasets demonstrated the effectiveness of our proposed method.
Supplementary Material
Organization of the Supplementary Material
  • Section A provides more details of the experimented perturbations and results of PR in CLIP backdoored by more attacks.
@@
  • Section C compares computational expenses between our method and CleanCLIP.
  • Section D contains more ablation studies, such as the visual encoder architecture, and the handcrafted prompt. It also contains the setting of experiments that ablates the FR loss.
  • Section E discusses the difference between adversarial training and the proposed RVPT.
A More Results of Perturbation Resistivity of More Backdoored CLIP
In the pilot experiment, we evaluate the model’s perturbation resistivity against 4 benign perturbations:
  • Gaussian Noise with a standard deviation of 0.001.
  • Gaussian Blur with a kernel size of 3 and variance of 5.
  • Color Jitter with brightness of 0.5 and hue of 0.3
  • Random Horizontal Flip.
The visualizations of these perturbations and trigger patterns can be found in Figure 7 and 10.
We evaluate the perturbation resistivity of four models:
  • ViT/B32 supervised exclusively on ImageNet.
  • ViT/B32 of OpenAI CLIP.
  • ViT/B32 of OpenAI CLIP tuned by VPT with 16 shots on ImageNet.
  • ViT/B32 of OpenAI CLIP tuned by RVPT with 16 shots on ImageNet.
The visual illustration for backdoored CLIP by BadNet and Blended are shown in Figure 8, 9 respectively. The results show that CLIP, even after VPT has generally much less PR value against various perturbations than traditionally supervised ViT, and RVPT can consistently increase its PR value across all scenarios.
Figure 7: Visualization of the perturbations of various experimented attacks.
Figure 8: Average PR values of CLIP backdoored by BadNet.
Figure 9: Average PR values of CLIP backdoored by Blended.
@@ B.1 Detailed Settings of Backdoor Attacks
In our experiments, we evaluate six prominent backdoor attack methods: BadNet [17], Blended [7], BadCLIP [31], ISSBA [28], WaNet [41], and TrojVQA [52]. Below, we provide a detailed description of each method:
@@
  • TrojVQA proposes a dual-key backdoor attack that sets triggers on both modalities. To evaluate RVPT under backdoored generative VLMs, we use TrojVQA as the backdoor attack and use the following settings. Firstly, we optimize the visual triggers for the CLIP visual backbone with the description “This is a yellow banana.”. Specifically, We set the patch size to 16$\times$16 and located the patch in the middle of the image. Then, we randomly corrupt the text for poisoned samples into sentences that describe the target class of banana. Moreover, we add the trigger word “SUDO ” in front of each sentence to make sure two keys for one backdoor. In the inference stage, the optimized trigger is added to the image, and “SUDO ” is added to the beginning of the prompt.
We implement the attacks following [3]. Specifically, we employ the AdamW optimizer with an initial learning rate of 1e-6 for all the backdoor attack methods. The learning rate follows a cosine decay schedule over five epochs, and we set the batch size to 128. The visualization of each trigger pattern is in Figure 10.
@@ B.2 Detailed Settings of Backdoor Defenses
CleanCLIP  CleanCLIP [3] provides a defense against backdoor attacks in multimodal contrastive learning by optimizing the integration of multimodal contrastive and unimodal self-supervised losses, relying on a limited amount of clean data. It is important to note that CleanCLIP utilizes ResNet-50 as the backbone of its visual encoder. In contrast, our work adopts the Vision Transformer (ViT-B/32) as the visual encoder, and we have accordingly adjusted the parameters to fit this architecture. Specifically, we randomly selected 250,000 image-text pairs from the CC3M dataset for fine-tuning. The learning rates were set to 5e-6 for BadNet, TrojVQA, Blended, and BadCLIP, while for WaNet and ISSBA on ImageNet-1K, they were set to 3e-6. We used a batch size of 64 and fine-tuned the models over 10 epochs. Notably, we did not focus solely on reducing the attack success rates by manipulating learning rates; instead, we ensured that the clean accuracy of the fine-tuned model was consistently maintained throughout.
@@ B.3 Detailed Settings of Datasets
C Computational Expenses
We conducted the main experiments on eight single NVIDIA 3090 GPUs using a half-precision data type and recorded one training process in Table 7. The results indicate that our method substantially reduces computational costs, enhancing defense efficiency.
CleanCLIP 3:53:02 17640 MB 126 M 250K
RVPT on ImageNet 45:05 3382 MB 0.34 M 16 K
RVPT on OxfordPets 2:19 1010 MB 0.34 M 0.6 K
@@ D More Experiments for Ablation Study
D.1 Ablation Study of the Handcrafted Prompt
@@
           “<CLS>”            0.09 (62.28)            0.03 (62.20)            3.72 (61.73)
           “###### <CLS>”            0.05 (62.18)            0.02 (61.96)            3.78 (61.37)
           “a photo of a <CLS>”            0.05 (62.76)            0.02 (62.36)            2.76 (61.81)
D.2 Ablation Study of the Visual Encoder Architecture
@@
D.3 Settings for Experiment of Ablation Study
@@
  • For the evaluation of open-vocabulary ASR, we tune the infected model whose target class is banana with the few-shot training samples on ImageNet, which excludes samples of the target class.
@@
E Discussion about the Difference between RVPT and Adversarial Training
As illustrated in 3, RVPT adopts the FR loss to adversarially repel features to impede the learning process, which partly shares some similarity with adversarial training (AT) [2]. While this approach shares some superficial similarities with adversarial training (AT) [2], there are fundamental differences between the two methodologies.
First, AT generates the predictive yet brittle features [21] via gradient-based attacks [16, 38], and then unlearns them with correct labels. (first pick bad features then unlearn) In contrast, RVPT first actively scrambles the feature representations and then learns the most predictive features. (first scramble features then pick good ones)
As a result, AT prevents the model from encoding the predictive, brittle features, while RVPT encourages the model to encode the in-dataset predictive features. In the training process of RVPT, there are no adversarial features unlearned. Therefore, it cannot defend against adversarial attacks. Table 10 compares the performance of RVPT and a representative adversarial training method of CLIP under adversarial attack. However, there are non-predictive features or out-of-dataset features unlearned, which ensures RVPT to defend against backdoor attacks.
Table 10: Accuracy of RVPT and adversarial Prompt Tuning [63] of adversarial samples on ImageNet. The attack is PGD-10, which maximizes the cross-entropy loss with a budge of 8/255.
No defense RVPT Adversarial Prompt Tuning
Accuracy 4.37 4.10 13.97